If a manufacturing company reacts 25 kilograms of 52.5% calcium nitrate ore with excess ammonium phosphate, how many grams of calcium phosphate are produced?
1
Expert's answer
2016-02-12T00:01:01-0500
3Ca(NO3)2->Ca3(PO4)2
m(CaNO3)=25*1000*0.525=13125 g
N(Ca(NO3)2)=80 moles
N(Ca3(PO4)2)= 80/3 = 26.67 moles; m(Ca3(PO4)2) = 8270 g